OpenStudy (anonymous):

help!!! solve the differential eqn: x''-4x=e^(-2t) given x(0)=0 and x'(0)=1

OpenStudy (turingtest):

what's the complimentary?

OpenStudy (anonymous):

compl= Ae^(2t)+Be^(-2t)

OpenStudy (turingtest):

so what is your guess for the particular?

OpenStudy (anonymous):

i tried x= Ae^(-2t) x'=-2Ae^(-2t) x''=4Ae^(-2t)

OpenStudy (turingtest):

you can't do that because we already have a term Ae^(-2t) in the complimentary, and we can't have any parts of the complimentary and the particular linearly dependent

OpenStudy (turingtest):

how do we normally deal with two linearly dependent solutions, like in a double root?

OpenStudy (anonymous):

ow... that's why... there's a rule? wow

OpenStudy (anonymous):

we add a "t" to C2

OpenStudy (turingtest):

yes, the rule is that no terms y1, y2,... ,yn or yp can be linearly dependent if they are linearly dependent we have to multiply something by the independent variable usually

OpenStudy (turingtest):

add a t to the guess for the particular in this case

OpenStudy (anonymous):

x= Ate^(-2t) x'=-2Ate^(-2t) x''=4Ate^(-2t)

OpenStudy (turingtest):

careful, you need to use the product rule in that derivative

OpenStudy (anonymous):

ow yes... wait

OpenStudy (anonymous):

x= Ate^(-2t) x'=Ae^(-2t)-2tAe^(-2t) x''=-2Ae^(-2t)-2Ae^(-2t)+4Ate^(-2)

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

:) i'm getting better at this A=-1/4?? Yp=(-1/4)te^(-2t)

OpenStudy (turingtest):

yes you are, that's the right particular :)

OpenStudy (anonymous):

now my initial conditions x(0)=0 gives A+B=0

OpenStudy (turingtest):

yes, and the other condition?

OpenStudy (anonymous):

A=-5/16 B=5/16

OpenStudy (turingtest):

yes, nice!

OpenStudy (turingtest):

actually I think you switched A and B, but besides that, nice job

OpenStudy (anonymous):

LOL... i just realized that... i couldn't find my x(1)... in one of the MCQs given... but i got it now.. thanks

OpenStudy (turingtest):

welcome!

OpenStudy (anonymous):

i actually learnt a lot from u since last night bernouli eqn particular solns... for a linear i use mx+c, quadratic ax^2+bx+c, one that has the same as my particular soln i add a t :)

OpenStudy (anonymous):

one last question before i go to bed :)